When a 3.80 g sample of liquid octane (C8H18) burns in a bomb calorimeter, the temperature of the calorimeter rises by 27.3 °C. The heat capacity of the calorimeter, measured in a separate experiment, is 6.18 kJ/°C. What is ∆E for the combustion for octane?
1
Expert's answer
2020-04-18T07:00:41-0400

Heat generated by 3.8 grams of octane = 27.3x6.18 =168.714 kJ

moles of octane in 3.8 grams = 3.8/114 =0.0333

heat generated by 0.0333 moles of octane is 168.714 kJ

heat generated by 1 moles of octane is "\frac{168.714}{0.0333}=5066.48 kJ"

this is called "\Delta E" of octane